# ]Meta FAIR Improving Factuality with Explicit Working Memory

### v2

@@ Abstract
Mingda Chen at
@@ 1 Introduction
Ewe can be seen as a more general framework that subsumes many existing approaches. For example, if there is no stopping in generation and if the memory contains only one unit (i.e., =1), then Ewe degenerates to the simple vanilla RAG. If Ewe pauses at the end of generation of every sentence and checks whether the new sentence contains any token with low probability as a proxy of factuality measure, then this particular instantiation, with one memory unit, is effectively FLARE. Notice that in a typical, more general configuration of Ewe, the memory module consists of multiple units. When the memory is refreshed, not all the units need to be updated. If some knowledge is still required, their original raw data (e.g., passages) will not be reprocessed to create the embeddings, saving some redundant computational cost at inference time. While conceptually simple, the working memory design in Ewe provides a more flexible and yet efficient way to incorporate various of types of external online information, as different forms of feedback are encoded in parallel and stored in memory units (e.g., see Figure 1). We notice that the design of leveraging working memory is also very related to some recently proposed methods for long-content models (e.g., Memory3 (Yang et al., 2024)). If our memory is used only for encoding the knowledge from the passages in our corpus, then this can be viewed as the whole corpus is used as the context, along with the prompt, as the input to the model. The key differences are that Ewe does not pre-encode every passage (although the KV caches of some frequently retrieved passages can certainly be precomputed in advance) and its memory can be dynamically updated as the generation progresses.
We demonstrate empirically that Ewe generates more factual responses without sacrificing the relevance to the input questions, using four fact-seeking long-form generation datasets. In general, with the feedback from online fact-checking and targeted retrieval, Ewe increases VeriScore (Song et al., 2024), the factuality metric we use, by 2 to 6 points absolute and has a similar helpfulness in terms of instruction following compared to the base model Llama-3.1. We also perform human evaluations to ensure that our enhancements in VeriScore effectively translate to factuality.
@@ 2.1 Iterative and Adaptive Retrieval Augmentation
Our work shares similarities with Self-RAG (Asai et al., 2024), particularly in the use of an auxiliary model to provide feedback. Unlike confidence measures based on token probability or attention score, Self-RAG fine-tunes the model to introspectively decide when to pause generation by outputting a special retrieve token. This triggers the retrieval of multiple passages, which are then used separately to generate candidate segments via standard RAG. Each segment is evaluated by a “critique” model for relevance, usefulness to the original prompt, and support from the retrieved passage. The critique model’s output determines whether a candidate segment is included in the final output.
@@ 3.1 Real-time Feedback
We pause at every timesteps to gather feedback from retrievers, and timesteps from fact-checkers. However, if no new sentence is generated, the feedback collection process will be skipped.
@@ 3.2 Refreshing Working Memories
The working memory in Ewe consists of memory units, where each unit is designed to store the representations of each feedback message of tokens. When updating working memories, we follow the first in, first out (FIFO) rule. Given refreshed text chunks of the same length from fact-checkers and retrievers, our model encodes them into the KV cache in parallel using the same positional IDs. Working memories in Ewe are stored as part of the language models’ context preceding the model’s own output text and prompts, allowing for flexible updates without reprocessing generated content. As shown in Figure 2, a separate embedding store is used for preserving these memories, which are then processed at each layer by concatenating them with the context. We then apply regular self-attention and aggregate the resulting hidden vectors using normalization terms from self-attention for each memory unit. Empirically, we find that adding hidden vectors produced by context only improves the fluency of long outputs, so we keep it in our model architectures. More formally,
@@ 4.1 Evaluation Datasets
We evaluated Ewe and baseline models using four fact-seeking long-form generation datasets: LongFact (Wei et al., 2024), Fava (Mishra et al., 2024), AlpacaFact (Dubois et al., 2023; Lin et al., 2024) and Biography (Min et al., 2023).
LongFact
Fava
AlpacaFact
Biography
@@ 4.2 Evaluation Metrics
We assess the quality of model responses to fact-seeking questions based on two key axes: factuality and helpfulness. For evaluating factuality, we considered multiple automatic metrics, such as FActScore (Min et al., 2023) and SAFE (Wei et al., 2024), but ultimately chose VeriScore (Song et al., 2024) as our primary evaluation metric. Although these metrics share a similar design that decomposes sentences into “atomic claims” and checks their support against an external knowledge source, VeriScore focuses on extracting more sensible verifiable claims and uses Google search snippets instead of Wikipedia as the knowledge source. As a result, VeriScore can be applied to responses on more diverse topics and is also more efficient, requiring fewer but more meaningful claims to be checked. We report the F1 score from VeriScore, which is the harmonic mean of the precision and recall of the claims. Following Song et al. (2024), we set the minimum number of facts required for a model’s response to achieve perfect recall as the median number of extracted claims per dataset111The median numbers of extracted facts for LongFact, Fava, AlpacaFact, Biography are 55, 49, 31, 43, respectively.. We also used their fine-tuned models for claim extraction and verification, provided in their package222https://github.com/Yixiao-Song/VeriScore.
To make sure that a model with a high factuality score does not simply give irrelevant but correct factual statements, we also need to check whether the response is helpful to the user. Following Lin et al. (2024), we use AlpacaEval (Dubois et al., 2024) to compare the target model and baseline model in terms of their instruction-following ability. For the responses to the same input prompt, a large language model is used as judge to determine which of the two is better333We used GPT-4o as the judge., and the win rate is thus used as a measure of helpfulness444We found that the length-controlled win rates in AlpacaEval could conflate hallucinations and length effects, and thus report the version without length normalization..
4.3 Baselines
We used instruction-tuned Llama-3.1 70B and 8B as the base models and compared Ewe with five baselines: base model only, retrieval augmentation (RA), Chain of verification (CoVe)555We adapted an implementation from https://github.com/ritun16/chain-of-verification, an iterative retrieval approach DRAGIN (Su et al., 2024)666We used the authors’ implementation https://github.com/oneal2000/DRAGIN, and a recently proposed semi-parametric decoding method Nest (Li et al., 2024). For base model only, Llama-3.1 or Llama-3.1, we simply gave the language model the prompt in the dataset and the instruction of requesting detailed information, without other additional information. With retrieval augmentation, we retrieved 10 passages using the input prompts as queries and then prepended the passages to the input777Using more than 10 passages does not provide significant benefits in our preliminary experiments, so we limit our retrieval to the top 10 passages.. Nest is a strong retrieval-based decoding algorithm. Following the original setup, we retrieved 100 passages to use as candidates. For CoVe, we employ the “factor+revise” method, which Dhuliawala et al. (2024) demonstrated to be the most effective. Additionally, we improve CoVe by integrating retrieved passages from our retrieval datastore during the verification step. This augmentation helps us establish a stronger and more comparable baseline method, considering that most other baseline methods also utilize retrieval. For all our experiments, the maximum generation step was set to 1024. Llama-3.1 is used as the baseline method for all AlpacaEval comparisons.
4.4 Results
@@ 5 Analysis
We provide some insights based on different ablations in this section. Llama-3.1 were used as the base model for all the experiments in this section.
@@ 5.1 Memory Configurations
In Figure 4, we examine the impact of varying lengths of memory units on factuality. Our results indicate that shorter memory units tend to increase precision but decrease recall, whereas longer memory units tend to decrease precision but increase recall. We hypothesize that this occurs because short memory units allow the attention mechanism to allocate weights more effectively to individual passages. In contrast, longer memory units combine multiple passages into a single unit, which may compel the attention to focus on less relevant passages when they are grouped with more relevant ones. However, longer memory units also provide advantages in terms of improved reasoning across documents, which appear to be beneficial for recall.
@@
5.2 Feedback Forms
@@
5.3 Model Confidence
Fixed intervals for refreshing working memory
As shown in Figure 6(a), when using a fixed retrieval interval, an smaller interval seems to perform better in general with a few spikes. This may be due to the fact that overly frequent retrieval can add irrelevant and conflicting information to the memory. On the other hand, when using a fixed verification interval, shorter interval always leads to better performance.
Model confidence for refreshing working memory
In practice, fixed retrieval and verification intervals may be unnecessary and lead to sub-optimal performance. We explore whether model-confidence can serve as a signal for refreshing working memory. Specifically, we compare two different metrics for model confidence: (1) Entropy: average entropy of generated tokens in a sentence, and (2) Min-prob: minimum probability of tokens in a sentence. A higher threshold for entropy results in less frequent memory update, and a higher threshold for min-prob results in more frequent memory update. Since external fact-checkers can be computationally expensive, we first examine if we can use model confidence as a signal for retrieval and regeneration, without using an auxiliary fact-checking model to provide feedback. As shown in Figure 6(b) and 6(c) (blue line), we observe empirically intermediate thresholds for retrieval perform well, leading to to better F1 when compared to the settings in Figure 6(a), where we use different fixed intervals for retrieval. With external fact-checkers, we investigate if we can use model confidence as a signal to trigger verification and regeneration to improve generation efficiency. In Figure 6(b) and 6(c) (green line), when chosen at an appropriate threshold, both entropy and min-prob can outperform the baseline (using fixed verification interval with the same memory configuration) despite with less frequent verification.
5.4 Knowledge from Retrieval
| Datastore | LongFact | Biography | AlpacaFact | Fava |
| Wiki | 67.9 | 46.1 | 55.5 | 52.5 |
| C4 | 70.8 | 44.6 | 53.7 | 53.3 |
| C4 + Wiki | 74.8 | 48.4 | 53.3 | 52.3 |
5.5 Human Evaluation
| RA | 0.61 |
| Ewe | 0.65 |
We conduct a human evaluation to determine if using VeriScore during inference impacts its alignment with human judgments. To achieve this, we randomly select 100 generation samples each from our retrieval-augmented baseline and Ewe, and recruit 6 human annotators. During sample selection, we categorize sentences into buckets based on the number of nonfactual claims predicted by VeriScore. We then randomly choose samples from each bucket to ensure an equal representation of sentences with varying numbers of nonfactual claims. Annotators are given a sentence from model responses along with the full response for context. They are instructed to classify the sentence as factual, nonfactual, or “unable to decide” if no claim is present. Additionally, they are encouraged to use online search engines if they are unsure about specific information. Our evaluation gathers 120 annotations. Since annotators might choose “unable to decide” for various reasons, such as the sentence being functional and not making any claims, we exclude these from score computation. We calculate Cohen’s Kappa to assess inter-annotator agreement between humans and VeriScore, excluding 25 “unable to decide” annotations. The results are presented in Table 4. Since a Cohen’s Kappa score above 0.61 is considered substantial and the scores do not decrease with Ewe, we conclude that using VeriScore during inference has indeed enhanced the factuality of Ewe.
6 Conclusion
We present Ewe, a novel system that incorporates a working memory mechanism during the generation process. Ewe pauses at given intervals and refreshes its working memory based on feedback from retrieval and fact-checking models, ensuring that the generated content remains accurate and relevant. By integrating this working memory into each attention layer of the Transformer architectures, Ewe can be easily adapted to various large language models. Our experiments demonstrate the effectiveness of Ewe by benchmarking it on 8B and 70B Llama-3.1 models, resulting in significant improvements in both factuality and helpfulness across four fact-seeking long-form generation datasets. Furthermore, our analysis reveals that updating the working memory with more relevant information at each timestep, allowing attention to focus on each passage, and utilizing high-quality retrieval datastores with extensive knowledge coverage are crucial factors for improving factuality of models.
Acknowledgments
We thank Victoria Lin, Barlas Oğuz, and Xilun Chen for their insightful feedback, and we are grateful to Maria Lomeli for her support in building the retrieval system.

### v3

# Improving Factuality with Explicit Working Memory

@@ Abstract
Improving Factuality with Explicit Working Memory
Mingda Chen Yang Li Karthik Padthe Rulin Shao Alicia Sun Luke Zettlemoyer Gargi Ghosh Wen-tau Yih Meta FAIR mingdachen@meta.com
@@ 1 Introduction
Ewe can be seen as a more general framework that subsumes many existing approaches. For example, if there is no stopping in generation and if the memory contains only one unit (i.e., =1), then Ewe degenerates to the simple vanilla RAG. If Ewe pauses at the end of generation of every sentence and checks whether the new sentence contains any token with low probability as a proxy of factuality measure, then this particular instantiation, with one memory unit, is effectively FLARE. Notice that in a typical, more general configuration of Ewe, the memory module consists of multiple units. When the memory is refreshed, not all the units need to be updated. If some knowledge is still required, their original raw data (e.g., passages) will not be reprocessed to create the embeddings, saving some redundant computational cost at inference time. While conceptually simple, the working memory design in Ewe provides a more flexible and yet efficient way to incorporate various types of external online information, as different forms of feedback are encoded in parallel and stored in memory units (e.g., see Figure 1). It is worth noting that Ewe is well-suited for real-time applications involving streaming user data, iterative retrieval, and real-time feedback. For example, when users input data through speech, where edits to previously entered information are less likely to occur, Ewe allows us to start gathering relevant information and drafting responses concurrently. We notice that the design of leveraging working memory is also very related to some recently proposed methods for long-content models (e.g., Memory3 (Yang et al., 2024)). If our memory is used only for encoding the knowledge from the passages in our corpus, then this can be viewed as the whole corpus is used as the context, along with the prompt, as the input to the model. The key differences are that Ewe does not pre-encode every passage (although the KV caches of some frequently retrieved passages can certainly be precomputed in advance) and its memory can be dynamically updated as the generation progresses.
We demonstrate empirically that Ewe generates more factual responses without sacrificing the relevance to the input questions, using four fact-seeking long-form generation datasets. In general, with the feedback from online fact-checking and targeted retrieval, Ewe increases VeriScore (Song et al., 2024), the factuality metric we use, by 2 to 6 points absolute and is equally helpful in terms of instruction following, compared to the base model Llama-3.1. We also perform human evaluations to ensure that our enhancements in VeriScore effectively translate to factuality.
@@ 2.1 Iterative and Adaptive Retrieval Augmentation
Our work shares similarities with Self-RAG (Asai et al., 2024), particularly in the use of an auxiliary model to provide feedback. Unlike confidence measures based on token probability or attention score, Self-RAG fine-tunes the model to introspectively decide when to pause generation by outputting a special retrieve token. This triggers the retrieval of multiple passages, which are then used separately to generate candidate segments via standard RAG. Each segment is evaluated by a “critique” model for relevance, usefulness to the original prompt, and support from the retrieved passage. The critique model’s output determines whether a candidate segment is included in the final output. Similarly, previous research has investigated various methods for incorporating external feedback during the generation process (Kang et al., 2023; Gao et al., 2023; Gou et al., 2024; Li and Flanigan, 2024).
@@ 3.1 Real-time Feedback
We pause at every timesteps to gather feedback from retrievers, and timesteps from fact-checkers. However, if no new sentence is generated, the feedback collection process will be skipped.111Ewe is not significantly slower than existing iterative retrieval methods like DRAGIN and FLARE. These methods also use an iterative generation process involving re-generation and retrieval. However, without working memories, existing iterative retrieval methods are slower than EWE when it comes to incorporating new retrieved information into the context.
@@ 3.2 Refreshing Working Memories
The working memory in Ewe consists of memory units, where each unit is designed to store the representations of each feedback message of tokens. When updating working memories, we follow the first in, first out (FIFO) rule.222In our preliminary experiments, we explored other possibilities of updating the memories, e.g., replacing the least relevant memories based on retrieval scores or model confidence scores (similar to what you mentioned). However, we found that these methods did not perform better than the FIFO approach. Due to the simplicity and effectiveness of FIFO, we decided to leave further exploration of these alternatives for future work. Given refreshed text chunks of the same length from fact-checkers and retrievers, our model encodes them into the KV cache in parallel using the same positional IDs. Working memories in Ewe are stored as part of the language models’ context preceding the model’s own output text and prompts, allowing for flexible updates without reprocessing generated content. As shown in Figure 2, a separate embedding store is used for preserving these memories, which are then processed at each layer by concatenating them with the context. We then apply regular self-attention and aggregate the resulting hidden vectors using normalization terms from self-attention for each memory unit. Empirically, we find that adding hidden vectors produced by context only improves the fluency of long outputs, so we keep it in our model architectures. More formally,
@@ 4.1 Evaluation Datasets
We evaluated Ewe and baseline models using four fact-seeking long-form generation datasets: LongFact (Wei et al., 2024; 250 prompts), Fava (Mishra et al., 2024; 141 prompts), AlpacaFact (Dubois et al., 2023; Lin et al., 2024; 241 prompts) and Biography (Min et al., 2023; 181 prompts). We include more detailed descriptions of each dataset in the appendix.
@@ 4.2 Evaluation Metrics
We assess the quality of model responses to fact-seeking questions based on two key axes: factuality and helpfulness. For evaluating factuality, we considered multiple automatic metrics, such as FActScore (Min et al., 2023) and SAFE (Wei et al., 2024), but ultimately chose VeriScore (Song et al., 2024) as our primary evaluation metric. Although these metrics share a similar design that decomposes sentences into “atomic claims” and checks their support against an external knowledge source, VeriScore focuses on extracting more sensible verifiable claims and uses Google search snippets instead of Wikipedia as the knowledge source. As a result, VeriScore can be applied to responses on more diverse topics and is also more efficient, requiring fewer but more meaningful claims to be checked. We report the F1 score from VeriScore, which is the harmonic mean of the precision and recall of the claims. Following Song et al. (2024), we set the minimum number of facts required for a model’s response to achieve perfect recall as the median number of extracted claims per dataset333The median numbers of extracted facts for LongFact, Fava, AlpacaFact, Biography are 55, 49, 31, 43, respectively.. We also used their fine-tuned models for claim extraction and verification, provided in their package444https://github.com/Yixiao-Song/VeriScore.
To make sure that a model with a high factuality score does not simply give irrelevant but correct factual statements, we also need to check whether the response is helpful to the user. Following Lin et al. (2024), we use AlpacaEval (Dubois et al., 2024) to compare the target model and baseline model in terms of their instruction-following ability. For the responses to the same input prompt, a large language model is used as judge to determine which of the two is better555We used GPT-4o as the judge., and the win rate is thus used as a measure of helpfulness666We found that the length-controlled win rates in AlpacaEval could conflate hallucinations and length effects, and thus report the version without length normalization..
@@
4.3 Baselines
We used instruction-tuned Llama-3.1 70B and 8B as the base models777We opt to use Llama-3.1 in two different sizes instead of exploring different architectures. This decision is based on the fact that Ewe is not tied to any specific model architecture. Consequently, it is more insightful to assess its generalization capabilities across various model sizes rather than across different architectures. and compared Ewe with five baselines: base model only, retrieval augmentation (RA), Chain of verification (CoVe)888We adapted an implementation from https://github.com/ritun16/chain-of-verification, an iterative retrieval approach DRAGIN (Su et al., 2024)999We used the authors’ implementation https://github.com/oneal2000/DRAGIN, and a recently proposed semi-parametric decoding method Nest (Li et al., 2024). For base model only, Llama-3.1 or Llama-3.1, we simply gave the language model the prompt in the dataset and the instruction of requesting detailed information, without other additional information. With retrieval augmentation, we retrieved 10 passages using the input prompts as queries and then prepended the passages to the input101010Using more than 10 passages does not provide significant benefits in our preliminary experiments, so we limit our retrieval to the top 10 passages.. Nest is a strong retrieval-based decoding algorithm. Following the original setup, we retrieved 100 passages to use as candidates. For CoVe, we employ the “factor+revise” method, which Dhuliawala et al. (2024) demonstrated to be the most effective. Additionally, we improve CoVe by integrating retrieved passages from our retrieval datastore during the verification step. This augmentation helps us establish a stronger and more comparable baseline method, considering that most other baseline methods also utilize retrieval. For all our experiments, the maximum generation step was set to 1024. We also use C4 and Wikipedia in MassiveDS as our retrieval datastore and Contriever as the retriever. Llama-3.1 is used as the baseline method for all AlpacaEval comparisons.
4.4 Results
@@ 5 Analysis
We provide some insights based on different ablations in this section. Llama-3.1 were used as the base model for all the experiments in this section. Due to space limits, we include analysis on model confidence, retrieval knowledge, and feedback forms for fact verifiers in the appendix.
@@ 5.1 Memory Configurations
In Figure 4, we examine the impact of varying lengths of memory units on factuality. Our results indicate that shorter memory units tend to increase precision but decrease recall, whereas longer memory units tend to decrease precision but increase recall. We hypothesize that this occurs because short memory units allow the attention mechanism to allocate weights more effectively to individual passages. In contrast, longer memory units combine multiple passages into a single unit, which may compel the attention to focus on less relevant passages when they are grouped with more relevant ones. However, longer memory units also provide advantages in terms of improved reasoning across documents, which is beneficial for recall.
@@
5.2 Human Evaluation
| RA | 0.61 |
| Ewe | 0.65 |
We conduct a human evaluation to determine if using VeriScore during inference impacts its alignment with human judgments. To achieve this, we randomly select 100 generation samples each from our retrieval-augmented baseline and Ewe, and recruit 6 human annotators. During sample selection, we categorize sentences into buckets based on the number of nonfactual claims predicted by VeriScore. We then randomly choose samples from each bucket to ensure an equal representation of sentences with varying numbers of nonfactual claims. Annotators are given a sentence from model responses along with the full response for context. They are instructed to classify the sentence as factual, nonfactual, or “unable to decide” if no claim is present. Additionally, they are encouraged to use online search engines if they are unsure about specific information. Detailed instructions are provided in the appendix. Our evaluation gathers 120 annotations. Since annotators might choose “unable to decide” for various reasons, such as the sentence being functional and not making any claims, we exclude these from score computation. We calculate Cohen’s Kappa to assess inter-annotator agreement between humans and VeriScore, excluding 25 “unable to decide” annotations. The results are presented in Table 2. Since a Cohen’s Kappa score above 0.61 is considered substantial and the scores do not decrease with Ewe, we conclude that using VeriScore during inference has indeed enhanced the factuality of Ewe.
6 Conclusion
We present Ewe, a novel system that incorporates a working memory mechanism during the generation process. Ewe pauses at given intervals and refreshes its working memory based on feedback from retrieval and fact-checking models, ensuring that the generated content remains accurate and relevant. Our experiments demonstrate the effectiveness of Ewe by benchmarking it on 8B and 70B Llama-3.1 models, resulting in significant improvements in both factuality and helpfulness across four fact-seeking long-form generation datasets. Furthermore, our analysis reveals that updating the working memory with more relevant information at each timestep, allowing attention to focus on each passage, and utilizing high-quality retrieval datastores with extensive knowledge coverage are crucial factors for improving factuality of models.
7 Limitations
Although Ewe can be applied to various types of data, our research has been limited to English text datasets with fact-seeking prompts. Its performance on other English benchmarks and multilingual datasets remains unclear. Additionally, we have restricted Ewe to only accept textual real-time feedback, and exploring multimodal feedback could be intriguing. Moreover, our human evaluation experiments have been conducted on a relatively small scale, and it would be beneficial to validate our findings in a larger-scale setting.
Appendix A Evaluation Datasets
LongFact
Fava
AlpacaFact
Biography
Appendix B Analysis on Model Confidence
Fixed intervals for refreshing working memory
As shown in Figure 9, when using a fixed retrieval interval, an smaller interval seems to perform better in general with a few spikes. This may be due to the fact that overly frequent retrieval can add irrelevant and conflicting information to the memory. On the other hand, when using a fixed verification interval, shorter interval always leads to better performance.
Model confidence for refreshing working memory
In practice, fixed retrieval and verification intervals may be unnecessary and lead to sub-optimal performance. We explore whether model-confidence can serve as a signal for refreshing working memory. Specifically, we compare two different metrics for model confidence: (1) Entropy: average entropy of generated tokens in a sentence, and (2) Min-prob: minimum probability of tokens in a sentence. A higher threshold for entropy results in less frequent memory update, and a higher threshold for min-prob results in more frequent memory update. Since external fact-checkers can be computationally expensive, we first examine if we can use model confidence as a signal for retrieval and regeneration, without using an auxiliary fact-checking model to provide feedback. As shown in Figure 9 and 9 (blue line), we observe empirically intermediate thresholds for retrieval perform well, leading to to better F1 when compared to the settings in Figure 9, where we use different fixed intervals for retrieval. With external fact-checkers, we investigate if we can use model confidence as a signal to trigger verification and regeneration to improve generation efficiency. In Figure 9 and 9 (green line), when chosen at an appropriate threshold, both entropy and min-prob can outperform the baseline (using fixed verification interval with the same memory configuration) despite with less frequent verification.
Appendix C Knowledge from Retrieval
| Datastore | LongFact | Biography | AlpacaFact | Fava |
| Wiki | 67.9 | 46.1 | 55.5 | 52.5 |
| C4 | 70.8 | 44.6 | 53.7 | 53.3 |
| C4 + Wiki | 74.8 | 48.4 | 53.3 | 52.3 |
Appendix D Feedback Forms
@@
Please refrain from including the following imprecise statements: (1) nonfactual claim1 (2) nonfactual claim2 …
@@
Appendix E Instructions for Human Annotators
Below is the instruction we provide to each annotator.
You will be given a sentence extracted from a model response, along with the full model response and the original prompt. Your task is to assess the factuality of the provided sentence. You may use Google search to help in your evaluation. A sentence should be considered as factual only if you can locate sources that corroborate all the claims made within it. If the sentence contains no claims, you may select ”unable to decide.”